OpenStudy (anonymous):

Find the second derivative of the function f(x) = log5 x^3.

OpenStudy (anonymous):

I know the first derivative is \[\frac{ 3 }{ xln5 }\] but then i'm stuck...

OpenStudy (anonymous):

If I were you I'd start by using the change of base formula to convert the logarithm to a quotient of natural logarithms, which are really easy to derive. OR Using what you just typed: 3/ln5 is a constant isn't it?

OpenStudy (anonymous):

What do you think now, in light of the rules you may have learned?

OpenStudy (anonymous):

Well I think I need to use the chain rule... but I don't know for sure... And I have no idea what the derivative of xln5 would be...

OpenStudy (anonymous):

Constant multiple rule. What's the derivative of 3/x?

OpenStudy (anonymous):

-3/x^2?

OpenStudy (anonymous):

Yes, and since ln5 is a constant, you can just insert it into the denominator. Maybe I made this difficult.\[\frac{ d }{ dx }(\frac{ 1 }{ x })=\frac{ -1 }{ x ^{2} }\]\[\frac{ d }{ dx }(\frac{ 3 }{ xln 5})=\frac{ 3 }{ \ln 5 }* \frac{ d }{ dx }(\frac{ 1 }{ x })\]because 3/ln5 is a constant. Check it on a calculator if you like

OpenStudy (anonymous):

hmmm... the answer i'm given is \[-\frac{ 3 }{ \ln5 } x^-2\]

OpenStudy (anonymous):

Think about it: What does a negative exponent do?

OpenStudy (anonymous):

oh pellet... makes it a fraction.

OpenStudy (anonymous):

Yep. Btw, are you at all interested in using the change of base formula on future problems?

OpenStudy (anonymous):

yes i am interested.

OpenStudy (anonymous):

i need to learn it... well i'm supposed to already know it. i'm a little behind in my class.

OpenStudy (anonymous):

That's all right. Math is something you can learn any time\[\log_{b}x=\frac{ \ln x }{ \ln b } \] or (treating ln b as a constant)\[=\frac{ 1 }{ \ln b }(\ln x)\] This saves you from learning a rule just for derivatives of logs in other bases. The derivatives of natural logs are really easy.

OpenStudy (anonymous):

If you need any more help, just mention me (that @ thing) or message me

OpenStudy (anonymous):

okay. thanks a lot for your help today.

OpenStudy (anonymous):

yw